OpenStudy (liv1234):

Use the binomial expression (p+q)n to calculate a binomial distribution with n=5 and p=0.3. @phi

OpenStudy (phi):

p= 0.3 and q = 1-p = 0.7 I think they want you to expand (p+q)^n using http://www.purplemath.com/modules/binomial.htm and then calculate the numerical value for each term (each value of n)

OpenStudy (liv1234):

Can you help me with that? That's where I'm really confused. @phi

OpenStudy (phi):

start by writing down the binomial expansion formula \[ (p+q)^n= \sum_{n=0}^n \left(\begin{matrix}n \\ k\end{matrix}\right)p^k\ (1-p)^{n-k}\]

OpenStudy (liv1234):

Would I plug in the numbers I already have into that formula?

OpenStudy (phi):

let me fix the index: \[ (p+q)^n= \sum_{k=0}^n \left(\begin{matrix}n \\ k\end{matrix}\right)p^k\ (1-p)^{n-k} \]

OpenStudy (phi):

there are going to be 6 terms, k=0, 1, ... , 5

OpenStudy (liv1234):

Okay, so what would be next?

OpenStudy (phi):

Let's do the first term, with k=0, n= 5, p= 0.3, q= 0.7

OpenStudy (phi):

it's easier to type if we use n Choose K, rather than the paren notation (it means the same thing)

OpenStudy (liv1234):

\[(0.3+0.7)^5 = \sum_{k=0}^{}\]

OpenStudy (liv1234):

Would it be something like that?

OpenStudy (phi):

"sort of" the binomial distribution has a "value" for each k and if we plot it, it will look like a gaussian if we have a large n. at each k, we want to get a number. If we plot a bar graph (each bar at a different k) we would see the distribution. They want us to figure out the values for each k.

OpenStudy (phi):

How do we figure out the values for each k (from 0 to 5) ? we use the "binomial theorem" to find each term.

OpenStudy (phi):

for example, for k=0 the first term is 5 C 0 0.3^0 0.7^5

OpenStudy (liv1234):

Okay, so next would be k=1?

OpenStudy (phi):

yes, give it a try.

OpenStudy (liv1234):

What was the second term?

OpenStudy (liv1234):

5 C 2 0.3^2 0.7^3?

OpenStudy (phi):

that is the 3rd term, with k=2

OpenStudy (liv1234):

5 C 1 0.3^1 0.7^4?

OpenStudy (phi):

yes

OpenStudy (liv1234):

What's next?

OpenStudy (phi):

do the same for k=3,4,5 also, give the numerical equivalent for each expression and then make a table, or a plot. For example, here is a plot of the binomial distribution with p=0.3 and n= 5

OpenStudy (liv1234):

5C3 .3^3 .7^2 5C4 .3^4 .7^1 5C5 .3^5

OpenStudy (phi):

notice for such a small n (n=5), it does not look like a bell-curve.

OpenStudy (liv1234):

What would we do next?

OpenStudy (phi):

can you evaluate each of those expressions ?

OpenStudy (liv1234):

Yes, but how would I do that?

OpenStudy (liv1234):

Can you help me please?

OpenStudy (phi):

n C k means n! / k! (n-k)! my calculator has a key for this operation.

OpenStudy (phi):

5C0 is by definition 1 also 5C5 is 1 5C1 is 5 also 5C4 is 5 that leaves 5C2 , 5C3

OpenStudy (liv1234):

Okay, and that means that goes into the calculator?

OpenStudy (phi):

let's do the first term. 5C0 0.3^0 0.7^5 what do you get for this ?

OpenStudy (liv1234):

I'm not sure, my calculator doesn't do that type of operation.

OpenStudy (phi):

you know 5C0 (I told you up above). it is 1 you know 0.3^0

OpenStudy (liv1234):

.16807

OpenStudy (phi):

yes, so that is the value for k=0 (or x=0 if we plot this ... see my graph) now the 2nd term 5C1 0.3^1 0.7^4

OpenStudy (liv1234):

.36015

OpenStudy (phi):

yes. btw, google knows how to do 5 choose 2 next term, for k=2. 5 C 2 0.3^2 0.7^3

OpenStudy (liv1234):

.3087

OpenStudy (liv1234):

Is that correct?

OpenStudy (phi):

yes. can you finish? they just want a table with those values

OpenStudy (liv1234):

I got the rest of the values of each term, what is the next step?

OpenStudy (phi):

that is the answer: a table with k and the corresponding value

OpenStudy (phi):

or perhaps use x: x Pr(x) 0 0.16807 1 0.36015 etc

OpenStudy (liv1234):

Thank you!

OpenStudy (phi):

and label it: binomial distribution for p=0.3 and n=5